OpenStudy (anonymous):

if sec = 4 then that means cos = 4 ??

Nnesha (nnesha):

what is the definition of sec for example tan = sin/cos sec =what ? :-0

OpenStudy (anonymous):

sec = 1/cos

Nnesha (nnesha):

yes right given cos = 4 so replace cos by 4 :-)

Nnesha (nnesha):

yes right given sec = 4 so replace 4 :-)

Nnesha (nnesha):

remember sec and cos are RECIPROCAL of each other

OpenStudy (anonymous):

so sec = 4 means 4/1 and cos = 1/4

Nnesha (nnesha):

yes right :-)

OpenStudy (anonymous):

Kk thanks.

Nnesha (nnesha):

np :-)